OpenStudy (anonymous):

Heres a simple x^3 one - i just forgot how to do them. f(x)=x^3 - 2x^2 - 13x - 10 as (x-a) (x-b) (x-c) [Simply the 3 factors for this equation]. (x+2) is one of them.

OpenStudy (anonymous):

could someone answer an show as (x+2(x-b)(x-c) [usually (x-a)(x-b)(x-c) but using the FACTOR THEOREM]

OpenStudy (anonymous):

to find the other factors your can either divide or think. you know it will be \[x^3 - 2x^2 - 13x - 10(x+2)(\text{ something })\] and you can figure out the something by

OpenStudy (anonymous):

i know thats the part im stuck at D:

OpenStudy (anonymous):

by division, by synthetic division, or by pure reason. you know the leading term must be x^2 and the constant must be -5 then figure out what the middle term has to be

OpenStudy (anonymous):

but you cannot use the factor theorem unless you know the zeros. if you know the zeros then you can use it to factor

OpenStudy (anonymous):

\[x^3 - 2x^2 - 13x - 10=(x+2)(x^2 + \text{ something } -5)\] we know the -5 because the constant is -10, and we know the x^2 because the leading term is x^3 now we can reason out what "something" has to be because -5 times x is -5x but you have -13x so you need -8 x, which comes by multiplying 2 by -4x

OpenStudy (anonymous):

I been using it by \[x ^{3} - 2x ^{2} - 13x - 10 = (ax ^{2} + bx +c)(x+2)+r\] \[x ^{3}+2x^{2}-13x-10=ax^{3} + 2ax^{2}+ bx^{2} +2bx +cx +2c + r\] (Comparing x^3) 1=a (Comparing x^2) 2=2a+2b 2=2+0 b=0 (Comparing x) -13=2b+c -13=0+c c=-13 (Comparing 0) -10=2c+r -10=-26+r r=16 so therefore \[x ^{3} - 2x ^{2} - 13x - 10 = (x ^{2} + 0x -13)(x+2)+16\] but from here i cant work it out >:/

OpenStudy (anonymous):

oh sorry for Comparing x^2 2=2a+b